# Dynamics of Agentic Loops in Large Language Models: A Geometric Theory of Trajectories From Semantic Contraction to Exploratory Divergence

### v2

@@ 2 Related Work
The question of implementing computational creativity in iterative systems provides the broader motivation for this work. Boden boden2009computermodels established the foundational framework for computational creativity, identifying three primary processes: exploration (searching conceptual spaces for new combinations), transformation (modifying the rules or structure of conceptual spaces) and combination (combining ideas from different domains). Colton and Wiggins colton2012computational formalized creativity through measurable dimensions: novelty (semantic or structural distance from existing examples), value (quality, coherence and usefulness) and surprise (unexpectedness relative to prior expectations). Zhao et al. zhao2024creativity recently demonstrated that large language models show creative behaviors across multiple tasks and domains.
@@
Madaan et al. madaan2023selfrefine introduced Self-Refine, demonstrating that large language models can iteratively critique and improve their own outputs through a feedback loop without additional training. Their approach achieved substantial improvements across diverse tasks including dialogue generation and mathematical reasoning. However, Self-Refine focused on empirical task performance rather than analyzing the geometric structure of how representations evolve through iterations or characterizing convergence properties in semantic space.
Wei et al. wei2022chain showed that chain-of-thought prompting enables large language models to perform complex reasoning by decomposing problems into intermediate steps. While this work established that prompting strategies significantly affect reasoning capabilities, it did not examine the stability or trajectory dynamics of iterative processes.
@@
Recent work has explored how language models can act as autonomous agents through iterative interaction with environments or tools. Yao et al. yao2022react proposed ReAct, which interleaves reasoning traces with task-specific actions, enabling models to interact with external knowledge sources. Shinn et al. shinn2023reflexion introduced Reflexion, where agents use verbal reinforcement learning to maintain episodic memory and improve decision-making across trials. Park et al. park2023generative demonstrated generative agents that simulate believable human behavior through architectures combining memory, reflection and planning.
@@
Understanding the geometric structure of embedding spaces is crucial for analyzing semantic trajectories. Ethayarajh ethayarajh2019contextual systematically studied contextualized representations from BERT, ELMo and GPT-2, demonstrating that embeddings become increasingly anisotropic in upper layers, with vectors concentrating in a narrow cone. This anisotropy causes high cosine similarities even between semantically distant items, distorting similarity-based measurements.
Wang and Isola wang2020understanding analyzed contrastive representation learning through the lens of alignment (bringing similar items together) and uniformity (distributing representations across the hypersphere). They showed that good representations balance these competing objectives, providing theoretical grounding for why pretrained embeddings exhibit systematic geometric biases.
For calibrating similarity metrics, Niculescu-Mizil and Caruana niculescu2005predicting demonstrated that isotonic regression effectively transforms classifier scores into well-calibrated probabilities. While originally developed for classification tasks, isotonic regression provides a principled non-parametric method for aligning model outputs with empirical distributions. We extend this approach to semantic similarity calibration.
@@
In contrast to existing work, which primarily focuses on static properties of embeddings or isolated transformations, the present work studies the dynamical behavior of iterative LLM transformations as geometric trajectories. Prior research on iterative refinement madaan2023selfrefine ; wei2022chain and agentic systems yao2022react ; shinn2023reflexion ; park2023generative has demonstrated impressive capabilities but treats each iteration as an independent step rather than as part of a continuous dynamical process. Similarly, embedding geometry research ethayarajh2019contextual ; wang2020understanding characterizes static distributional properties without examining how representations evolve under repeated transformations. Our work bridges these threads by treating agentic loops as discrete dynamical systems in semantic space, enabling the characterization of convergence, divergence and attractor structures through calibrated geometric measurements.
@@
While computational creativity frameworks boden2009computermodels ; colton2012computational motivate the need to measure trajectory-based metrics, prior work has not developed the foundational geometric measurement infrastructure required to rigorously analyze iterative transformation dynamics. Self-refinement approaches madaan2023selfrefine demonstrate empirical task improvements but do not characterize how representations evolve geometrically through iterations or analyze convergence versus divergence behaviors. Agentic frameworks yao2022react ; shinn2023reflexion ; park2023generative establish recursive LLM capabilities but lack quantitative measurement of trajectory dynamics in semantic space. Although embedding anisotropy is well-documented ethayarajh2019contextual and theoretical frameworks for representation quality exist wang2020understanding , no existing work provides calibrated, human-aligned similarity metrics for rigorous trajectory analysis in iterative transformations.
Our contribution builds the foundational geometric measurement infrastructure required for analyzing agentic loop dynamics, providing the tools necessary to eventually address computational creativity questions systematically. We address these gaps through three main innovations. First, we develop a geometric framework operating in the representation space treating iterative transformations as discrete dynamical systems. This framework captures measurable trajectories, clusters, attractors and dynamical regimes that quantify how semantic representations evolve under recursive application. These geometric constructs open the path to provide to measure Colton & Wiggins’ creativity criteria. Second, we introduce isotonic calibration that aligns embedding similarities with human semantic judgments extending calibration methods niculescu2005predicting to overcome anisotropy ethayarajh2019contextual and enable precise geometric measurements using metrics that reflect human-perceived semantic distances. Third, we provide systematic empirical characterization of dynamical regimes. We demonstrate that prompt design directly determines whether agentic loops exhibit convergence, divergence or complex behaviors, as revealed by measurable geometric indicators in calibrated semantic space. This establishes the foundation for engineering Boden’s creative processes through controlled trajectory dynamics.
@@ 3.4 High-Dimensional Similarity Distribution
On a standard high-dimensional unit sphere $\mathbb{S}^{d-1}$, two random vectors $u,v$ drawn uniformly from the surface tend to be nearly orthogonal as the dimension increases. Formally, let $X$ be uniformly distributed on $\mathbb{S}^{d-1}$ and let $f(X)=\langle X,v\rangle$ for a fixed $v\in\mathbb{S}^{d-1}$. By Lévy’s concentration inequality (see vershynin2018high for a modern treatment), for any $t>0$,
@@
However, the semantic embedding space induced by a pretrained language model such as paraphrase-mpnet-base-v2 diverges from this theoretical isotropy. Because the model was trained on non-uniform pairs of semantically related sentences, the embeddings are not uniformly distributed over the hypersphere but rather cluster around a dominant mean direction in $\mathbb{R}^{d}$. This induces a strong anisotropy of the embedding manifold, where most vectors share a large positive cosine with the mean embedding vector. As formalized by wang2020understanding , good representation learning requires balancing two properties: alignment (similar points should be close) and uniformity (embeddings should be distributed uniformly on the unit sphere). The observed concentration violates the uniformity property, indicating that the pretrained model prioritizes alignment over uniform distribution. This trade-off is typical of contrastively trained embeddings. Empirically, the distribution of $s^{(m)}(e_{1},e_{2})$ is concentrated around $0.8$, in stark contrast with the isotropic baseline centered at 0.
@@ 3.7 Visualization of Similarity Distributions
Silverman’s rule silverman1986density selects the optimal bandwidth $h$, balancing smoothness and resolution. We compute separate densities for human scores and model similarities, $\hat{f}_{H}(x)$ and $\hat{f}_{S}(x)$ and plot them over $[0,1]$. Alignment between both curves indicates good agreement, while systematic shifts reveal under- or over-confidence in the model’s similarity predictions. Additionally, both visualizations include a vertical green line marking the high-confidence similarity threshold $\tau_{\text{HCS}}^{(m)}$ (defined in §3.6), which serves as a data-driven decision boundary for determining when two sentences can be considered semantically similar with high confidence.
@@ 8.4 Application: Creativity as Objective
The agentic geometric cognition framework introduced in Section 8.3 raises a natural application question: what if the objective that the agent targets is to maximize creativity? This section explores how computational creativity, formalized by Boden boden2009computermodels as exploration-transformation-combination processes and by Colton and Wiggins colton2012computational as novelty-value-surprise dimensions, can be operationalized as a specific objective within the geometry of agentic cognition framework.
@@
This operationalization enables measurable creativity control. Boden’s exploratory creativity can be triggered within agentic loop iterations by introducing high-temperature sampling and role-based divergent prompts zhao2024creativity (for instance, adopting perspectives such as an artist, a child or a poet), which induce high dispersion and novelty generation. However, unconstrained exploration risks semantic drift and loss of value. To balance creativity, subsequent iterations can act as consolidators, employing role-based contractive prompts (e.g., scientist, engineer, critic) that reduce dispersion and restore coherence, maximizing value while preserving novel content.

### v3

@@ 2 Related Work
The question of implementing computational creativity in iterative systems provides the broader motivation for this work. Boden Boden (2009) established the foundational framework for computational creativity, identifying three primary processes: exploration (searching conceptual spaces for new combinations), transformation (modifying the rules or structure of conceptual spaces) and combination (combining ideas from different domains). Colton and Wiggins Colton and Wiggins (2012) formalized creativity through measurable dimensions: novelty (semantic or structural distance from existing examples), value (quality, coherence and usefulness) and surprise (unexpectedness relative to prior expectations). Zhao et al. Zhao et al. (2024) recently demonstrated that large language models show creative behaviors across multiple tasks and domains.
@@
Madaan et al. Madaan et al. (2023) introduced Self-Refine, demonstrating that large language models can iteratively critique and improve their own outputs through a feedback loop without additional training. Their approach achieved substantial improvements across diverse tasks including dialogue generation and mathematical reasoning. However, Self-Refine focused on empirical task performance rather than analyzing the geometric structure of how representations evolve through iterations or characterizing convergence properties in semantic space.
Wei et al. Wei et al. (2022) showed that chain-of-thought prompting enables large language models to perform complex reasoning by decomposing problems into intermediate steps. While this work established that prompting strategies significantly affect reasoning capabilities, it did not examine the stability or trajectory dynamics of iterative processes.
@@
Recent work has explored how language models can act as autonomous agents through iterative interaction with environments or tools. Yao et al. Yao et al. (2023) proposed ReAct, which interleaves reasoning traces with task-specific actions, enabling models to interact with external knowledge sources. Shinn et al. Shinn et al. (2023) introduced Reflexion, where agents use verbal reinforcement learning to maintain episodic memory and improve decision-making across trials. Park et al. Park et al. (2023) demonstrated generative agents that simulate believable human behavior through architectures combining memory, reflection and planning.
@@
Understanding the geometric structure of embedding spaces is crucial for analyzing semantic trajectories. Ethayarajh Ethayarajh (2019) systematically studied contextualized representations from BERT, ELMo and GPT-2, demonstrating that embeddings become increasingly anisotropic in upper layers, with vectors concentrating in a narrow cone. This anisotropy causes high cosine similarities even between semantically distant items, distorting similarity-based measurements.
Wang and Isola Wang and Isola (2020) analyzed contrastive representation learning through the lens of alignment (bringing similar items together) and uniformity (distributing representations across the hypersphere). They showed that good representations balance these competing objectives, providing theoretical grounding for why pretrained embeddings exhibit systematic geometric biases.
For calibrating similarity metrics, Niculescu-Mizil and Caruana Niculescu-Mizil and Caruana (2005) demonstrated that isotonic regression effectively transforms classifier scores into well-calibrated probabilities. While originally developed for classification tasks, isotonic regression provides a principled non-parametric method for aligning model outputs with empirical distributions. We extend this approach to semantic similarity calibration.
@@
In contrast to existing work, which primarily focuses on static properties of embeddings or isolated transformations, the present work studies the dynamical behavior of iterative LLM transformations as geometric trajectories. Prior research on iterative refinement Madaan et al. (2023); Wei et al. (2022) and agentic systems Yao et al. (2023); Shinn et al. (2023); Park et al. (2023) has demonstrated impressive capabilities but treats each iteration as an independent step rather than as part of a continuous dynamical process. Similarly, embedding geometry research Ethayarajh (2019); Wang and Isola (2020) characterizes static distributional properties without examining how representations evolve under repeated transformations. Our work bridges these threads by treating agentic loops as discrete dynamical systems in semantic space, enabling the characterization of convergence, divergence and attractor structures through calibrated geometric measurements.
@@
While computational creativity frameworks Boden (2009); Colton and Wiggins (2012) motivate the need to measure trajectory-based metrics, prior work has not developed the foundational geometric measurement infrastructure required to rigorously analyze iterative transformation dynamics. Self-refinement approaches Madaan et al. (2023) demonstrate empirical task improvements but do not characterize how representations evolve geometrically through iterations or analyze convergence versus divergence behaviors. Agentic frameworks Yao et al. (2023); Shinn et al. (2023); Park et al. (2023) establish recursive LLM capabilities but lack quantitative measurement of trajectory dynamics in semantic space. Although embedding anisotropy is well-documented Ethayarajh (2019) and theoretical frameworks for representation quality exist Wang and Isola (2020), no existing work provides calibrated, human-aligned similarity metrics for rigorous trajectory analysis in iterative transformations.
Our contribution builds the foundational geometric measurement infrastructure required for analyzing agentic loop dynamics, providing the tools necessary to eventually address computational creativity questions systematically. We address these gaps through three main innovations. First, we develop a geometric framework operating in the representation space treating iterative transformations as discrete dynamical systems. This framework captures measurable trajectories, clusters, attractors and dynamical regimes that quantify how semantic representations evolve under recursive application. These geometric constructs open the path to provide to measure Colton & Wiggins’ creativity criteria. Second, we introduce isotonic calibration that aligns embedding similarities with human semantic judgments extending calibration methods Niculescu-Mizil and Caruana (2005) to overcome anisotropy Ethayarajh (2019) and enable precise geometric measurements using metrics that reflect human-perceived semantic distances. Third, we provide systematic empirical characterization of dynamical regimes. We demonstrate that prompt design directly determines whether agentic loops exhibit convergence, divergence or complex behaviors, as revealed by measurable geometric indicators in calibrated semantic space. This establishes the foundation for engineering Boden’s creative processes through controlled trajectory dynamics.
@@ 3.4 High-Dimensional Similarity Distribution
On a standard high-dimensional unit sphere $\mathbb{S}^{d-1}$, two random vectors $u,v$ drawn uniformly from the surface tend to be nearly orthogonal as the dimension increases. Formally, let $X$ be uniformly distributed on $\mathbb{S}^{d-1}$ and let $f(X)=\langle X,v\rangle$ for a fixed $v\in\mathbb{S}^{d-1}$. By Lévy’s concentration inequality (see Vershynin (2018) for a modern treatment), for any $t>0$,
@@
However, the semantic embedding space induced by a pretrained language model such as paraphrase-mpnet-base-v2 diverges from this theoretical isotropy. Because the model was trained on non-uniform pairs of semantically related sentences, the embeddings are not uniformly distributed over the hypersphere but rather cluster around a dominant mean direction in $\mathbb{R}^{d}$. This induces a strong anisotropy of the embedding manifold, where most vectors share a large positive cosine with the mean embedding vector. As formalized by Wang and Isola (2020), good representation learning requires balancing two properties: alignment (similar points should be close) and uniformity (embeddings should be distributed uniformly on the unit sphere). The observed concentration violates the uniformity property, indicating that the pretrained model prioritizes alignment over uniform distribution. This trade-off is typical of contrastively trained embeddings. Empirically, the distribution of $s^{(m)}(e_{1},e_{2})$ is concentrated around $0.8$, in stark contrast with the isotropic baseline centered at 0.
@@ 3.7 Visualization of Similarity Distributions
Silverman’s rule Silverman (1986) selects the optimal bandwidth $h$, balancing smoothness and resolution. We compute separate densities for human scores and model similarities, $\hat{f}_{H}(x)$ and $\hat{f}_{S}(x)$ and plot them over $[0,1]$. Alignment between both curves indicates good agreement, while systematic shifts reveal under- or over-confidence in the model’s similarity predictions. Additionally, both visualizations include a vertical green line marking the high-confidence similarity threshold $\tau_{\text{HCS}}^{(m)}$ (defined in §3.6), which serves as a data-driven decision boundary for determining when two sentences can be considered semantically similar with high confidence.
@@ 8.4 Application: Creativity as Objective
The agentic geometric cognition framework introduced in Section 8.3 raises a natural application question: what if the objective that the agent targets is to maximize creativity? This section explores how computational creativity, formalized by Boden Boden (2009) as exploration-transformation-combination processes and by Colton and Wiggins Colton and Wiggins (2012) as novelty-value-surprise dimensions, can be operationalized as a specific objective within the geometry of agentic cognition framework.
@@
This operationalization enables measurable creativity control. Boden’s exploratory creativity can be triggered within agentic loop iterations by introducing high-temperature sampling and role-based divergent prompts Zhao et al. (2024) (for instance, adopting perspectives such as an artist, a child or a poet), which induce high dispersion and novelty generation. However, unconstrained exploration risks semantic drift and loss of value. To balance creativity, subsequent iterations can act as consolidators, employing role-based contractive prompts (e.g., scientist, engineer, critic) that reduce dispersion and restore coherence, maximizing value while preserving novel content.